OpenStudy (aryana_maria2323):

Find f '(x) for f of x equals the square root of the quantity sine of 8 times x. a. f prime of x equals the quotient of the cosine of 8 times x and 2 times the square root of the quantity sine of 8 times x b. f prime of x equals the quotient of 4 times the cosine of 8 times x and the square root of the quantity sine of 8 times x c. f prime of x equals the quotient of 4 and the square root of the quantity sine of 8 times x d. f'(x) = sin(4x)

OpenStudy (aryana_maria2323):

@123AB456C @tkhunny

zepdrix (zepdrix):

\[\large\rm f(x)=\sqrt{\sin8x}\]

zepdrix (zepdrix):

Hey there :) This is a handy derivative that you'll want to put to memory,\[\large\rm \frac{d}{dx}\sqrt x=\frac{1}{2\sqrt x}\]Derivative of square root is `one over two square roots`. You can do your 1/2 power if you prefer. Square root shows up so often though, that it's worth remembering :)

OpenStudy (aryana_maria2323):

The derivative for f(x)=\sqrt{\sin 8x}\] is \[4\csc(8x)12\cos(8x)\]

zepdrix (zepdrix):

Hmm where did your square root go? :O It vanished! Let's apply our square root derivative,\[\large\rm f(x)=\sqrt{\sin8x}\]so then,\[\large\rm f'(x)=\frac{1}{2\sqrt{\sin8x}}\cdot(\sin8x)'\]And then of course we have chain rule on the outside.

zepdrix (zepdrix):

Your solution doesn't look far off, just curious about a few things.

OpenStudy (aryana_maria2323):

d/dxf(g(x))=f′(g(x))g′(x) isn't this the way to find a derivative using the chain rule?

zepdrix (zepdrix):

Yes. Does something I said not seem to match up with that? :o

OpenStudy (aryana_maria2323):

No I was just making sure I was doing it right.

zepdrix (zepdrix):

So our derivative of the inner function, our g(x)=sin8x, gives us\[\large\rm f'(x)=\frac{1}{2\sqrt{\sin8x}}\cdot(8\cos8x)\]

zepdrix (zepdrix):

So I'm not quite sure where your 12 is coming from, hmm

OpenStudy (aryana_maria2323):

okay after apply the chain rule to the problem you gave me I got \[2\csc(8x)\frac{ 1 }{ 2} \cos(8x)\]

OpenStudy (aryana_maria2323):

you still there @zepdrix

OpenStudy (aryana_maria2323):

Can you please help? @ Nnesha @triciaal I think @zepdrix left.

OpenStudy (triciaal):

where did you get 12 from? how did you simplify 1/(squareroot Sin 8x) to get

zepdrix (zepdrix):

Ah sorry bout that D:

OpenStudy (triciaal):

review what zepdrix has

OpenStudy (aryana_maria2323):

it wasn't supposed to be 12. I think when I typed it in here I forgot the /. Its supposed to be 1/2

zepdrix (zepdrix):

Oh oh I see, you intended for that 1/2 to be the exponent. Ok this is making more sense :)

OpenStudy (triciaal):

ok and he is back

OpenStudy (aryana_maria2323):

ohh no problem @zepdrix. I have been having computer problems and I didn't know if is was from my end.

zepdrix (zepdrix):

Keep in mind that the 2 should stay in the denominator,\[\large\rm \frac{1}{2(\sin8x)^{1/2}}\quad=\quad \frac{(\csc8x)^{1/2}}{2}\]

OpenStudy (aryana_maria2323):

Yah I am new to Open study so I am still trying to figure out how to work everthing

OpenStudy (aryana_maria2323):

now do I find the derivative of (csc8x) 1/2 2

zepdrix (zepdrix):

Again, I would strongly recommend you try using the square root derivative :)) But if you like your 1/2 powers, \[\large\rm f(x)=(\sin8x)^{1/2}\]just proceed carefully,\[\large\rm f'(x)=\frac12(\sin8x)^{-1/2}\cdot(\sin8x)'\]\[\large\rm f'(x)=\frac12(\sin8x)^{-1/2}\cdot(8\cos8x)\]

zepdrix (zepdrix):

Notice with your definition of chain rule, d/dx f(g(x)) = f'(g(x))*g'(x) See how the f'( ) contains the `original inside function`, the inner function doesn't change when you're applying this differentiation process. Chain rule tells you to instead make a copy of the inner function, and multiply by its derivative.

OpenStudy (aryana_maria2323):

Okay so the derivative of the question using the chain rule is. \[4\csc(8x)^\frac{ 1 }{ 2 }\cos(8x)\]

OpenStudy (aryana_maria2323):

I am not sure how to use the square root derivative. I was only taught to use the chain rule derivative

zepdrix (zepdrix):

I'm not sure where this 4 is coming from... So from this step,\[\large\rm f'(x)=\frac12(\sin8x)^{-1/2}\cdot(\sin8x)'\]Yes, you can apply your trig identity to the sine,\[\large\rm f'(x)=\frac12(\csc8x)^{1/2}\cdot(\sin8x)'\]So you have a 1/2 in front, right? And we'll need to fix up our sine derivative at the end as well. One small step in chain rule that you're missing :)

OpenStudy (aryana_maria2323):

wait why does your 1st problem have a -1/2 and you 2nd problem have a +1/2

zepdrix (zepdrix):

Applying one of our exponent rules gives us this,\[\large\rm (\sin8x)^{-1/2}\quad=\quad\frac{1}{(\sin8x)^{1/2}}\]And then trig identity from there. This is the exponent rule I'm thinking of:\[\large\rm x^{-a}=\frac{1}{x^a}\]

OpenStudy (aryana_maria2323):

Ohh okay so now I need to find the derivative of your problem with the +1/2 correct?

zepdrix (zepdrix):

Ah sorry, I mean, yes :) you need to find the derivative of the thing that has the ' on it.

OpenStudy (aryana_maria2323):

they both do lol

zepdrix (zepdrix):

From this step,\[\large\rm f'(x)=\frac12(\csc8x)^{1/2}\cdot(\sin8x)'\]we need to find derivative of sin8x, that is our chain rule step. Your cos8x was very close.

zepdrix (zepdrix):

But it turns out that you have to chain again! :)\[\large\rm (\sin8x)'\quad=\quad (\cos8x)\cdot(8x)'\]

OpenStudy (aryana_maria2323):

okay so apply the chain rule to \[(\cos 8x) \times (8x)\]

zepdrix (zepdrix):

it's not just 8x on the outside, it's the derivative of 8x, don't forget the ' in that step! :)

zepdrix (zepdrix):

I hope you understand this notation ok, when I leave a tick mark on something, it's just to "set up" the next step. It's telling us that we still need to take a derivative.

zepdrix (zepdrix):

\[\large\rm (\sin8x)'\quad=\quad (\cos8x)\cdot(8x)'\]Derivative of 8x? :)\[\large\rm \frac{d}{dx}8x=?\]

OpenStudy (aryana_maria2323):

8

zepdrix (zepdrix):

\[\large\rm (\sin8x)'\quad=\quad (\cos8x)\cdot(8)\]Good good good.

zepdrix (zepdrix):

\[\large\rm f'(x)=\frac12(\csc8x)^{1/2}\cdot(\sin8x)'\]\[\large\rm f'(x)=\frac12(\csc8x)^{1/2}\cdot(\cos8x)\cdot(8x)'\]\[\large\rm f'(x)=\frac12(\csc8x)^{1/2}\cdot(\cos8x)\cdot(8)\]Bring the 8 and 1/2 together as a final step. Multiply them.

zepdrix (zepdrix):

Hmm, I didn't realize you had multiply choice options, woops. Notice that `all of your options` involve sin8x, not csc8x. So we probably should not have used that trig property.\[\large\rm f'(x)=\frac12(\sin8x)^{-1/2}\cdot(\cos8x)\cdot(8)\]

OpenStudy (aryana_maria2323):

okay so do you want me to derivative that problem?

zepdrix (zepdrix):

No derivative is all done :) We did the hard part. Now is just simplification.

OpenStudy (aryana_maria2323):

Okay so what would be the steps? I am trying to follow put I am kindof having a hard time. Could you draw the steps on how to simplify

zepdrix (zepdrix):

\[\large\rm f'(x)=\frac12(\sin8x)^{-1/2}\cdot(\cos8x)\cdot(8)\]Let's reapply that exponent step,\[\large\rm f'(x)=\frac12\frac{1}{(\sin8x)^{1/2}}\cdot(\cos8x)\cdot(8)\]Multiplication is `commutative`, so we can multiply in any order. Let's bring the 8 to the front, and divide it by the 2,\[\large\rm f'(x)=\frac82\frac{1}{(\sin8x)^{1/2}}\cdot(\cos8x)\]Gives us,\[\large\rm f'(x)=4\frac{1}{(\sin8x)^{1/2}}\cdot(\cos8x)\]

zepdrix (zepdrix):

And then as a final step, maybe put the stuff into the numerator,\[\large\rm f'(x)=\frac{4(\cos8x)}{(\sin8x)^{1/2}}\]

zepdrix (zepdrix):

Oh maybe another good step would be to rewrite your 1/2 power as a square root again :)\[\large\rm f'(x)=\frac{4(\cos8x)}{\sqrt{\sin8x}}\]

OpenStudy (aryana_maria2323):

oh okay now that makes sense. I was wondering what you were supposed to do with the 1/2 power. But now it makes sense.

zepdrix (zepdrix):

Try to commit a lot of time to understanding/practicing the chain rule. It is much more complex than the other derivative rules. Blessings to you!

OpenStudy (aryana_maria2323):

Thank you so very much for all the help. I got the answer CORRECT!!